OpenStudy (anonymous):

A man has a 90-foot ladder and leans the end of it against an exterior window of a house. The base of the ladder is 45 feet from the house. Caculate from the nearest foot how high it is from window to the ground

OpenStudy (dangerousjesse):

So you have a right triangle 

OpenStudy (dangerousjesse):

\(a^2 + b^2 = c^2\) a=45 c=90

OpenStudy (dangerousjesse):

Solve for b and you have your answer :)

OpenStudy (anonymous):

ok thanks this was very helpful

OpenStudy (dangerousjesse):

No problem

OpenStudy (anonymous):

one more question do i add or mutiply for b

OpenStudy (anonymous):

She/he is currently offline

OpenStudy (anonymous):

And you add for B

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

sure